# THeGCN: Temporal Heterophilic Graph Convolutional Network

### v1

@@ Abstract
Introduction
 [^1] Preliminary work using AAAI template.
Graph Neural Network (GNN) (Kipf and Welling 2016; Veličković et al. 2017; Hamilton, Ying, and Leskovec 2017) has exhibited great power in a large number of graph learning tasks, such as node classification (Kipf and Welling 2016), link prediction (Zhang and Chen 2018), recommender systems (Wang et al. 2022) and many more. The homophily assumption, i.e., the connected node pairs tend to share the same labels, is widely adopted by many existing GNNs. Beyond that, most GNNs are designed for static graphs, whose topology and attributes are fixed once the graph is constructed.
@@ Introduction
Recently, graphs with more complex topology and attributes, such as heterophilic graphs and temporal graphs, have attracted increasing research attention. In the spatial domain, it is found that mainstream GNNs have difficulties handling the edge heterophily issue, which refers to the phenomenon that connected nodes may have distinct labels/classes in real-world graphs (Bo et al. 2021; Luan et al. 2021). Examples of edge homophily and edge heterophily are shown in Figure 1. Numerous GNNs (Li et al. 2022; Du et al. 2022; Zheng et al. 2022) have been proposed to solve the edge heterophily issue. For example, H2GCN (Zhu et al. 2020) samples distant neighbors for message-passing. FAGCN (Bo et al. 2021) and GPRGNN (Chien et al. 2020) use adaptive low/high-pass filters to make connected node pairs share similar/disparate embeddings. On the other hand, in the temporal domain, the real-world graphs keep evolving over time due to the ever-emerging events. To model such event-based continuous graphs, TGAT (Xu et al. 2020) and TGN (Rossi et al. 2020) transfer the homophily assumption from static GNNs to temporal graphs and utilize the attention technique for message-passing. To summarize, as shown in Table 1, SOTA has been addressing these two lines of work almost in parallel:
@@
Homophilic GCN (Kipf and Welling 2016) GAT (Veličković et al. 2017) SGC (Wu et al. 2019) TGN (Rossi et al. 2020) TGAT (Xu et al. 2020) JODIE (Kumar, Zhang, and Leskovec 2019)
Heterophilic FAGCN (Bo et al. 2021) GPRGNN (Chien et al. 2020) HOG-GCN (Wang and Zhang 2022) THeGCN (This paper)
@@
Preliminaries and Problem Definition
@@
The layer-wise message-passing and aggregation of GCN (Kipf and Welling 2016) is given by
@@
Edge heterophily. Edge heterophily describes to what extent edges tend to link nodes with disparate labels. The edge heterophily measurement  (Zhu et al. 2020) for static graphs is defined as (Zhu et al. 2020): $H(\mathcal{G})=\frac{\sum_{i,j,\mathbf{A}[i,j]=1}\langle y_{v_{i}}\neq y_{v_{j}}\rangle}{\sum_{i,j}\mathbf{A}[i,j]}\in[0,1],$ where $\langle x\rangle=1$ if $x$ is true and 00 otherwise. A graph is more homophilic for $H(\mathcal{G})$ closer to 00 or more heterophilic for $H(\mathcal{G})$ closer to $1$.
@@
Temporal Edge Heterophily
@@
where $y^{t_{2}}_{v_{j}}$ is the label of $v_{j}$ at time $t_{2}$ and $\{e^{t}_{v_{i},v_{j}}\}$ is the set of events/interactions occurring between nodes $v_{i}$ and $v_{j}$ during time interval $[t_{1},t_{2})$. From Eq. (2), we can observe that the target node is $v_{j}$ at time $t_{2}$ and $H(t_{1},t_{2},v_{j})$ measures the ratio of interactions/edges which link nodes with different labels during $[t_{1},t_{2})$. The relation between temporal and static edge heterophily measurements are: firstly, the time $t_{1}$ and $t_{2}$ are variables of $H(t_{1},t_{2},v_{j})$, which indicates that its value is non-stationary in different time intervals, even for the same node; secondly, while the label of node $v_{i}$ is fixed in the static edge heterophily measurement ($y_{v_{i}}$), it can change with the time in the temporal edge heterophily measurement ($y^{t}_{v_{i}}$) in Eq. (2); thirdly, the existing static edge heterophily measurement can be recovered as a special case of the temporal edge heterophily measurement if the node labels are fixed with time and the time interval is set as $[0,\infty)$.
Key Idea
The key idea of THeGCN is utilizing the low/high-pass signal filtering to address the co-existence of the edge heterophily issue and the temporal heterophily issue. For one thing, the low/high-pass filtering is a widely used technique in static heterophilic graph neural networks (Bo et al. 2021; Yan et al. 2024; Du et al. 2022). Specifically, the low-pass filtering targets at making embeddings of connected nodes become similar (e.g., $\mathbf{I}+\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ in GCN (Kipf and Welling 2016)) to capture the homophilic information of edges in the graph, while the high-pass filtering aims to make embeddings of connected nodes more distinguishable (e.g., $\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$), which is able to reflect the heterophilic information. Through adaptively adjusting the weights of the low/high-pass filters, these static heterophilic graph neural networks can well solve the edge heterophily issue (SMP block). For another thing, according to the definition of temporal edge heterophily measurement in Eq. (2), the proposed THeGCN should also take the temporal heterophily issue into consideration and encode the temporal information in the module (TMP block).
Algorithm Details
@@
Sampler. For node $v_{j}$ at time $t^{\prime}$, we first set the time interval as $[t_{0},t^{\prime})$, where the choice of the starting point $t_{0}$ is a hyper-parameter. Then, we adopt the parallel sampler in TGL (Zhou et al. 2022a) ($\mathtt{Sampler}(v^{t^{\prime}}_{j},h_{\textrm{max}},[t_{0},t^{\prime}),N_{\textrm{max}})$) to randomly sample a maximal $N_{\textrm{max}}$ number of interactions/edges $e^{t}_{v_{i},v_{j}}$ within $h_{\textrm{max}}$ hops of $v_{j}$, where $t\in[t_{0},t^{\prime})$  [^2] If the number of interactions/edges within $h_{\textrm{max}}$ hops of $v_{j}$ is smaller than $N_{\textrm{max}}$, all interactions/edges are sampled.. All events/interactions involved in the described process form the interaction set $\mathcal{S}_{v_{j}}^{t^{\prime}}=\{e^{t}_{v_{i},v_{j}}\}$ for $v^{t^{\prime}}_{j}$. For each interaction $e^{t}_{v_{i},v_{j}}$ in $\mathcal{S}_{v_{j}}^{t^{\prime}}$, we leverage the following time encoder to encode the time information, which has been validated to be effective in (Cong et al. 2022):
@@
Aggregator. With the sampled interaction set $\mathcal{S}_{v_{j}}^{t^{\prime}}$, the aggregator of THeGCN conducts the message-passing operation to encode all related information. The aggregator is composed of two connected blocks: the temporal message-passing (TMP) block and the static message-passing (SMP) block  [^3] If the model does not have the SMP block, $l=0$, which is equivalent to one-layer TMP block.. The TMP block collects information from node features $\mathbf{x}_{v_{i}^{t}}$ and $\mathbf{x}_{v_{j}^{t}}$ of node $v_{i}$ and $v_{j}$ at $t$, the attribute $\mathbf{m}_{e_{v_{i},v_{j}}^{t}}$  [^4] The attribute of the interaction is optional. of interaction $e_{v_{i},v_{j}}^{t}$, and time encoding $E(t^{\prime}-t)$. A multi-layer perceptron (MLP) (Gardner and Dorling 1998) is used to produce the weight $p_{v^{t}_{i},v^{t}_{j}}^{(0)}$ of the low-pass filter for the temporal interaction $e_{v_{i},v_{j}}^{t}$ happening to $v_{j}$ at $t^{\prime}$:
@@
Complexity Analysis & Discussion.
Complexity analysis. Here we give a brief time complexity analysis for the proposed THeGCN. The time complexity of the sampler in each epoch is $\mathcal{O}(|\mathcal{E}|)$ according to (Zhou et al. 2022a), where $\mathcal{E}$ is the interaction set in the event-based continuous graph $\mathcal{T}$. For the aggregator, one message passing layer in either the TMP or the SMP block has time complexity $\mathcal{O}(|\mathcal{E}|\times f)$, where $f$ is the output dimension of the corresponding layer.
Discussion. THeGCN can be viewed as a generalization of some existing, well-established static heterophilic graph neural networks and temporal graph neural networks. We illustrate this connection with some examples from these two categories. For static heterophilic graph neural networks, if we remove the temporal information and the interaction information from Eq. (4) and Eq. (5), THeGCN will degrade to FAGCN (Bo et al. 2021) and the gating block of GBK-GNN (Du et al. 2022). For temporal graph neural networks, if we exclude the process of learning the low-pass/high-pass weights and conduct the message-passing mechanism with equal weights, THeGCN will become the node encoder in GraphMixer (Cong et al. 2022). If the low/high-pass weight in THeGCN is limited to single non-negative attention weight, THeGCN becomes similar to TGAT (Xu et al. 2020). We provide ablation studies on the single non-negative attention weight in the experiment section.
Experiment
@@
Experiment Setup
Datasets. We use 5 datasets for evaluation, including 3 traffic datasets: PEMSBA, PEMSLA and PEMSSD, where labels of nodes change with time, 1 social network dataset Reddit (Fan, Yao, and Joe-Wong 2021) and 1 biological dataset Brain, where labels of nodes do not change with time. The detailed description of these datasets are attached in Appendix due to page limit. All the statistics are presented in Table 3. In addition, a metric named temporal changing ratio is proposed. This metric measures the ratio of nodes whose labels have changed. We can see that over 95% of nodes have changed the labels during this period, which also matches the motivation of this work. The labels of nodes in Reddit and Brain are constructed by (Fan, Yao, and Joe-Wong 2021) and remain the same during the whole time period, which is reflected in Table 3 that the temporal changing ratios are 0s. We randomly split labeled nodes  [^7] In PEMSBA/PEMSLA/PEMSSD, since labels for the same node change with time, the same node with different labels at different times are viewed as two labeled nodes. In Reddit and Brain, the same node only appears once in the training/validation/test set. in every dataset into 60/20/20% for training, validation, and testing.
@@
Baselines and metrics. We compare our model with 10 baseline methods, which can be divided into two groups: methods in the first group are temporal graph neural networks, including TGN (Rossi et al. 2020), JODIE (Kumar, Zhang, and Leskovec 2019), APAN (Wang et al. 2021), DySAT (Sankar et al. 2020) and TGAT (Xu et al. 2020). Baselines in the second group are representative static heterophilic/homophilic graph neural networks: GCN (Kipf and Welling 2016), APPNP (Klicpera, Bojchevski, and Günnemann 2018), FAGCN (Bo et al. 2021), GPRGNN (Chien et al. 2020) and GAT (Veličković et al. 2017). We adopt node classification accuracy as the metric for all methods.
@@
Related Works
Dynamic graph neural network (GNN). According to the types of dynamic graphs, dynamic graph neural networks (GNNs) can be divided into discrete-time dynamic GNNs (Goyal et al. 2018; Wang et al. 2020), and event-based continuous GNNs (a.k.a. temporal GNNs) (Nguyen et al. 2018; Rossi et al. 2020; Trivedi et al. 2019; Cong et al. 2022; Zhou et al. 2022a). Earlier dynamic graph neural networks (Goyal et al. 2018; Wang et al. 2020; You, Du, and Leskovec 2022; Bai et al. 2020) view the dynamic graph neural network as static graph snapshots, whose information is relatively coarse and thus is not the focus of our paper. Recently, a large number of temporal GNNs emerge to capture the information in event-based continuous graphs. To name a few, to capture the temporal dynamics, temporal point process is formulated in DyRep (Trivedi et al. 2019). Temporal motifs are adopted in CAW (Wang et al. 2020) to model the dynamics of the graph. TGAT (Xu et al. 2020) transfers the attention mechanism from GAT (Veličković et al. 2017) in static graphs to event-based continuous graphs. TGN (Rossi et al. 2020) is designed to memorize long term dependencies efficiently and TGL (Zhou et al. 2022a) optimizes the sampling process of existing event-based continuous graphs. GraphMixer (Cong et al. 2022) simplifies the complex structures of existing event-based continuous GNNs. However, all of the above event-based continuous GNNs are developed based on the temporal homophily assumption. The latest work GRETO  (Zhou et al. 2022b) notices the temporal heterophily issue, but it is a discrete-time dynamic GNN, which can not be applied to event-based continuous graphs. To our best knowledge, THeGCN is the first work on temporal heterophily in event-based continuous graphs.
Node classification on static heterophilic graphs. Most existing graph neural networks (GNNs) (Kipf and Welling 2016; Veličković et al. 2017; Hamilton, Ying, and Leskovec 2017) are developed based on the homophily assumption, referring to the phenomenon that connected node pairs tend to share the same labels. Recently, researchers pay more attention to heterophilic graphs, where connected node pairs tend to have disparate labels. From the spatial aspect, H2GCN (Zhu et al. 2020) and Geom-GCN (Pei et al. 2020) adopt a large neighborhood field for message-passing. From the spectral aspect, FAGCN (Bo et al. 2021) and GPRGNN (Chien et al. 2020) adaptively integrate high/low frequency signals with trainable parameters. In addition, HOG-GCN (Wang and Zhang 2022) and CPGNN (Zhu et al. 2021) propose an alternative message-passing mechanism to handle the homophily and the heterophily. Recent progresses in this domain include ACM-GCN (Luan et al. 2021, 2022), LINKX (Lim et al. 2021), BernNet (He et al. 2021), GloGNN (Li et al. 2022), and GBKGNN (Du et al. 2022). (Zheng et al. 2022) has conducted a comprehensive survey on this topic.
Conclusion and Limitation

### v2

@@ Abstract
1 Introduction
 [^1] Preliminary work using IJCAI template.
Graph Neural Network (GNN) Kipf and Welling (2016); Veličković et al. (2017); Hamilton et al. (2017) has exhibited great power in a large number of graph learning tasks, such as node classification Kipf and Welling (2016), link prediction Zhang and Chen (2018), recommender systems Wang et al. (2022) and many more. The homophily assumption, i.e., the connected node pairs tend to share the same labels, is widely adopted by many existing GNNs. Beyond that, most GNNs are designed for static graphs, whose topology and attributes are fixed once the graph is constructed.
@@ 1 Introduction
Recently, graphs with more complex topology and attributes, such as heterophilic graphs and temporal graphs, have attracted increasing research attention. In the spatial domain, it is found that mainstream GNNs have difficulties handling the edge heterophily issue, which refers to the phenomenon that connected nodes may have distinct labels/classes in real-world graphs Bo et al. (2021); Luan et al. (2021). Examples of edge homophily and edge heterophily are shown in Figure 1. Numerous GNNs Li et al. (2022); Du et al. (2022); Zheng et al. (2022) have been proposed to solve the edge heterophily issue. For example, H2GCN Zhu et al. (2020) samples distant neighbors for message-passing. FAGCN Bo et al. (2021) and GPRGNN Chien et al. (2020) use adaptive low/high-pass filters to make connected node pairs share similar/disparate embeddings. On the other hand, in the temporal domain, the real-world graphs keep evolving over time due to the ever-emerging events. To model such event-based continuous graphs, TGAT Xu et al. (2020) and TGN Rossi et al. (2020) transfer the homophily assumption from static GNNs to temporal graphs and utilize the attention technique for message-passing. To summarize, as shown in Table 1, SOTA has been addressing these two lines of work almost in parallel:
@@
Homophilic GCN Kipf and Welling (2016) GAT Veličković et al. (2017) SGC Wu et al. (2019) TGN Rossi et al. (2020) TGAT Xu et al. (2020) JODIE Kumar et al. (2019)
Heterophilic FAGCN Bo et al. (2021) GPRGNN Chien et al. (2020) HOG-GCN Wang and Zhang (2022) THeGCN (This paper)
@@
2 Preliminaries and Problem Definition
@@
The layer-wise message-passing and aggregation of GCN Kipf and Welling (2016) is given by
@@
Edge heterophily. Edge heterophily describes to what extent edges tend to link nodes with disparate labels. The edge heterophily measurement  Zhu et al. (2020) for static graphs is defined as Zhu et al. (2020): $H(\mathcal{G})=\frac{\sum_{i,j,\mathbf{A}[i,j]=1}\langle y_{v_{i}}\neq y_{v_{j}}\rangle}{\sum_{i,j}\mathbf{A}[i,j]}\in[0,1],$ where $\langle x\rangle=1$ if $x$ is true and 00 otherwise. A graph is more homophilic for $H(\mathcal{G})$ closer to 00 or more heterophilic for $H(\mathcal{G})$ closer to $1$.
@@ 3 Model
3.1 Temporal Edge Heterophily
@@
where $y^{t_{2}}_{v_{j}}$ is the label of $v_{j}$ at time $t_{2}$, $v_{i}$ is the node has directed edge connected to $v_{j}$, and $\{e^{t}_{v_{i},v_{j}}\}$ is the set of events/interactions occurring between nodes $v_{i}$ and $v_{j}$ during time interval $[t_{1},t_{2})$. From Eq. (2), we can observe that the target node is $v_{j}$ at time $t_{2}$ and $H(t_{1},t_{2},v_{j})$ measures the ratio of interactions/edges which link nodes with different labels during $[t_{1},t_{2})$. The relation between temporal and static edge heterophily measurements are: firstly, the time $t_{1}$ and $t_{2}$ are variables of $H(t_{1},t_{2},v_{j})$, which indicates that its value is non-stationary in different time intervals, even for the same node; secondly, while the label of node $v_{i}$ is fixed in the static edge heterophily measurement ($y_{v_{i}}$), it can change with the time in the temporal edge heterophily measurement ($y^{t}_{v_{i}}$) in Eq. (2); thirdly, the existing static edge heterophily measurement can be recovered as a special case of the temporal edge heterophily measurement if the node labels are fixed with time and the time interval is set as $[0,\infty)$.
3.2 Key Idea
The key idea of THeGCN is utilizing the low/high-pass signal filtering to address the co-existence of the edge heterophily issue and the temporal heterophily issue. For one thing, the low/high-pass filtering is a widely used technique in static heterophilic graph neural networks Bo et al. (2021); Yan et al. (2024); Du et al. (2022). Specifically, the low-pass filtering targets at making embeddings of connected nodes become similar (e.g., $\mathbf{I}+\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ in GCN Kipf and Welling (2016)) to capture the homophilic information of edges in the graph, while the high-pass filtering aims to make embeddings of connected nodes more distinguishable (e.g., $\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$), which is able to reflect the heterophilic information. Through adaptively adjusting the weights of the low/high-pass filters, these static heterophilic graph neural networks can well solve the edge heterophily issue (SMP block). For another thing, according to the definition of temporal edge heterophily measurement in Eq. (2), the proposed THeGCN should also take the temporal heterophily issue into consideration and encode the temporal information in the module (TMP block).
3.3 Algorithm Details
@@
Sampler. For node $v_{j}$ at time $t^{\prime}$, we first set the time interval as $[t_{0},t^{\prime})$, where the choice of the starting point $t_{0}$ is a hyper-parameter. Then, we adopt the parallel sampler in TGL Zhou et al. (2022a) ($\mathtt{Sampler}(v^{t^{\prime}}_{j},h_{\textrm{max}},[t_{0},t^{\prime}),N_{\textrm{max}})$) to randomly sample a maximal $N_{\textrm{max}}$ number of interactions/edges $e^{t}_{v_{i},v_{j}}$ within $h_{\textrm{max}}$ hops of $v_{j}$, where $t\in[t_{0},t^{\prime})$  [^2] If the number of interactions/edges within $h_{\textrm{max}}$ hops of $v_{j}$ is smaller than $N_{\textrm{max}}$, all interactions/edges are sampled.. All events/interactions involved in the described process form the interaction set $\mathcal{S}_{v_{j}}^{t^{\prime}}=\{e^{t}_{v_{i},v_{j}}\}$ for $v^{t^{\prime}}_{j}$. For each interaction $e^{t}_{v_{i},v_{j}}$ in $\mathcal{S}_{v_{j}}^{t^{\prime}}$, we leverage the following time encoder to encode the time information, which has been validated to be effective in Cong et al. (2022):
@@
Aggregator. With the sampled interaction set $\mathcal{S}_{v_{j}}^{t^{\prime}}$, the aggregator of THeGCN conducts the message-passing operation to encode all related information. The aggregator is composed of two connected blocks: the temporal message-passing (TMP) block and the static message-passing (SMP) block  [^3] If the model does not have the SMP block, $l=0$, which is equivalent to one-layer TMP block.. The TMP block collects information from node features $\mathbf{x}_{v_{i}^{t}}$ and $\mathbf{x}_{v_{j}^{t}}$ of node $v_{i}$ and $v_{j}$ at $t$, the attribute $\mathbf{m}_{e_{v_{i},v_{j}}^{t}}$  [^4] The attribute of the interaction is optional. of interaction $e_{v_{i},v_{j}}^{t}$, and time encoding $E(t^{\prime}-t)$. A multi-layer perceptron (MLP) Gardner and Dorling (1998) is used to produce the weight $p_{v^{t}_{i},v^{t}_{j}}^{(0)}$ of the low-pass filter for the temporal interaction $e_{v_{i},v_{j}}^{t}$ happening to $v_{j}$ at $t^{\prime}$:
@@
3.4 Complexity Analysis & Discussion.
Complexity analysis. Here we give a brief time complexity analysis for the proposed THeGCN. The time complexity of the sampler in each epoch is $\mathcal{O}(|\mathcal{E}|)$ according to Zhou et al. (2022a), where $\mathcal{E}$ is the interaction set in the event-based continuous graph $\mathcal{T}$. For the aggregator, one message passing layer in either the TMP or the SMP block has time complexity $\mathcal{O}(|\mathcal{E}|\times f)$, where $f$ is the output dimension of the corresponding layer.
Discussion. THeGCN can be viewed as a generalization of some existing, well-established static heterophilic graph neural networks and temporal graph neural networks. We illustrate this connection with some examples from these two categories. For static heterophilic graph neural networks, if we remove the temporal information and the interaction information from Eq. (4) and Eq. (5), THeGCN will degrade to FAGCN Bo et al. (2021) and the gating block of GBK-GNN Du et al. (2022). For temporal graph neural networks, if we exclude the process of learning the low-pass/high-pass weights and conduct the message-passing mechanism with equal weights, THeGCN will become the node encoder in GraphMixer Cong et al. (2022). If the low/high-pass weight in THeGCN is limited to single non-negative attention weight, THeGCN becomes similar to TGAT Xu et al. (2020). We provide ablation studies on the single non-negative attention weight in the experiment section.
4 Experiment
@@
4.1 Experiment Setup
Datasets. We use 5 datasets for evaluation, including 3 traffic datasets: PEMSBA, PEMSLA and PEMSSD, where labels of nodes change with time, 1 social network dataset Reddit Fan et al. (2021) and 1 biological dataset Brain, where labels of nodes do not change with time. The detailed description of these datasets are attached in Appendix due to page limit. All the statistics are presented in Table 3. In addition, a metric named temporal changing ratio is proposed. This metric measures the ratio of nodes whose labels have changed. We can see that over 95% of nodes have changed the labels during this period, which also matches the motivation of this work. The labels of nodes in Reddit and Brain are constructed by Fan et al. (2021) and remain the same during the whole time period, which is reflected in Table 3 that the temporal changing ratios are 0s. We randomly split labeled nodes  [^7] In PEMSBA/PEMSLA/PEMSSD, since labels for the same node change with time, the same node with different labels at different times are viewed as two labeled nodes. In Reddit and Brain, the same node only appears once in the training/validation/test set. in every dataset into 60/20/20% for training, validation, and testing.
@@ 4.1 Experiment Setup
Baselines and metrics. We compare our model with 10 baseline methods, which can be divided into two groups: methods in the first group are temporal graph neural networks, including TGN Rossi et al. (2020), JODIE Kumar et al. (2019), APAN Wang et al. (2021), DySAT Sankar et al. (2020) and TGAT Xu et al. (2020). Baselines in the second group are representative static heterophilic/homophilic graph neural networks: GCN Kipf and Welling (2016), APPNP Klicpera et al. (2018), FAGCN Bo et al. (2021), GPRGNN Chien et al. (2020) and GAT Veličković et al. (2017). We adopt node classification accuracy as the metric for all methods.
@@ 4.2 Results and Analysis
5 Related Works
Dynamic graph neural network (GNN). According to the types of dynamic graphs, dynamic graph neural networks (GNNs) can be divided into discrete-time dynamic GNNs Goyal et al. (2018); Wang et al. (2020), and event-based continuous GNNs (a.k.a. temporal GNNs) Nguyen et al. (2018); Rossi et al. (2020); Trivedi et al. (2019); Cong et al. (2022); Zhou et al. (2022a). Earlier dynamic graph neural networks Goyal et al. (2018); Wang et al. (2020); You et al. (2022); Bai et al. (2020) view the dynamic graph neural network as static graph snapshots, whose information is relatively coarse and thus is not the focus of our paper. Recently, a large number of temporal GNNs emerge to capture the information in event-based continuous graphs. To name a few, to capture the temporal dynamics, temporal point process is formulated in DyRep Trivedi et al. (2019). Temporal motifs are adopted in CAW Wang et al. (2020) to model the dynamics of the graph. TGAT Xu et al. (2020) transfers the attention mechanism from GAT Veličković et al. (2017) in static graphs to event-based continuous graphs. TGN Rossi et al. (2020) is designed to memorize long term dependencies efficiently and TGL Zhou et al. (2022a) optimizes the sampling process of existing event-based continuous graphs. GraphMixer Cong et al. (2022) simplifies the complex structures of existing event-based continuous GNNs. However, all of the above event-based continuous GNNs are developed based on the temporal homophily assumption. The latest work GRETO  Zhou et al. (2022b) notices the temporal heterophily issue, but it is a discrete-time dynamic GNN, which can not be applied to event-based continuous graphs. To our best knowledge, THeGCN is the first work on temporal heterophily in event-based continuous graphs.
Node classification on static heterophilic graphs. Most existing graph neural networks (GNNs) Kipf and Welling (2016); Veličković et al. (2017); Hamilton et al. (2017) are developed based on the homophily assumption, referring to the phenomenon that connected node pairs tend to share the same labels. Recently, researchers pay more attention to heterophilic graphs, where connected node pairs tend to have disparate labels. From the spatial aspect, H2GCN Zhu et al. (2020) and Geom-GCN Pei et al. (2020) adopt a large neighborhood field for message-passing. From the spectral aspect, FAGCN Bo et al. (2021) and GPRGNN Chien et al. (2020) adaptively integrate high/low frequency signals with trainable parameters. In addition, HOG-GCN Wang and Zhang (2022) and CPGNN Zhu et al. (2021) propose an alternative message-passing mechanism to handle the homophily and the heterophily. Recent progresses in this domain include ACM-GCN Luan et al. (2021, 2022), LINKX Lim et al. (2021), BernNet He et al. (2021), GloGNN Li et al. (2022), and GBKGNN Du et al. (2022). Zheng et al. (2022) has conducted a comprehensive survey on this topic.
6 Conclusion and Limitation